OpenStudy (anonymous):

Simplify the expression. 2(3y - 7) - 5y(2 - y)

OpenStudy (ajprincess):

distribute 2(3y-7)nd -5y(2-y) 2(3y-7)=2*3y-2*7=6y-14 -5y(2-y)=-5y*2-5y*-y=-10y+5y^2 6y-14-10y+5y^2 Combine the like terms. Does that help?

OpenStudy (anonymous):

not really

OpenStudy (ajprincess):

well where r u stuck?

OpenStudy (ajprincess):

@keeno plz tell me which part of my answer nt clear to u